16) San Diego company has the following information available at the end of the fiscal year: Finished goods inventory, January 1, 2015 12,000 units Finished goods inventory, December 31, 2015 14,000 units Work in process inventory, January 1, 2015 10,000 units Work in process inventory, December 31, 2015 11,000 units Raw materials inventory, January 1, 2015 1,000 units Raw materials inventory, December 31, 2015 5,000 units Actual fixed overhead cost rate $2.05 per unit Actual variable overhead cost rate $3.10 per unit Budgeted fixed overhead cost rate $2.00 per unit Budgeted variable overhead cost rate $3.00 per unit Assume operating income under absorption costing is $100,000. What is the difference in operating income between absorption costing and variable costing?
Answer #1
Difference in the operating income is due to the difference in fixed overheads included in the finished goods inventory at the beginning and the end when the absorption costing is used.
Fixed overheads in beginning FG inventory = 12000*2 = $        24,000
Fixed overheads in ending FG inventory = 14000*2 = $        28,000
Net Fixed overhead deferred = 28000-24000 = $          4,000
Absorption costing income will be more by $4000
Hence, the difference between the two methods = $          4,000